# An Objective Improvement Approach

### v4

@@ 2. Preliminaries
The value of a minimiser strategy $\sigma$, denoted by $\mathsf{val}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{val}_{\sigma}:v\mapsto\max_{\sigma_{\max}\in\Sigma_{\max}}\mathsf{out}(\rho(v,\sigma,\sigma_{\max}))\ ,$
The value of a maximiser strategy $\sigma$, denoted by $\mathsf{val}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{val}_{\sigma}:v\mapsto\min_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma))\ .$
@@
A positional maximiser (resp. minimiser) strategy $\sigma$ is said to be optimal if and only if $\mathsf{val}_{\mathcal{G}}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\mathcal{G}}(\sigma(v))$ holds for all maximiser (resp. minimiser) vertices.
Likewise, we define the value of a pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, denoted $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:V\rightarrow\mathbb{R}$, as $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:v\mapsto\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))$.
@@
If both of these strategies are optimal, we call $\sigma$ a joint co-optimal strategy. This is the case if, and only if, $\mathsf{val}_{\mathcal{G}}=\mathsf{val}_{\sigma}$ holds [^1] If $\mathsf{val}_{\sigma}=\mathsf{val}_{\mathcal{G}}$ holds, then, for every vertex $v$ and for $v^{\prime}=\sigma(v)$, $\mathsf{val}_{\sigma}(v)=w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}_{\sigma}(v^{\prime})$ holds. At the same time, for all maximiser vertices $v$ and edges $(v,v^{\prime\prime})$, it then holds that $\mathsf{val}_{\sigma}(v)\leq w_{(v,v^{\prime\prime})}+\lambda_{(v,v^{\prime\prime})}\mathsf{val}_{\sigma}(v^{\prime\prime})$, while, for all minimiser vertices $v$ and edges $(v,v^{\prime\prime})$, $\mathsf{val}_{\sigma}(v)\geq w_{(v,v^{\prime\prime})}+\lambda_{(v,v^{\prime\prime})}\mathsf{val}_{\sigma}(v^{\prime\prime})$ holds. This implies that the minimiser and maximiser strategies constituting $\sigma$ are co-optimal..
@@
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution [^3] If $\sigma$ is not co-optimal, then there is a maximiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}_{\sigma}(v)<w_{e}+\lambda_{e}\mathsf{val}_{\sigma}(v^{\prime})$ or there is a minimiser vertex $v$ and an edge $e=(v,v^{\prime})$ such that $\mathsf{val}_{\sigma}(v)>w_{e}+\lambda_{e}\mathsf{val}_{\sigma}(v^{\prime})$. In both cases, the inequation for this particular edge $e$ is violated..
@@ 2.1. Simplex method
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that minimises (or maximises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Given a game $\mathcal{G}$, we can define an objective function represented by the vector c, the input matrix $A$, and the vector $b$, such that the solution $x$ to the linear programming instance corresponds to $\mathsf{val}_{\mathcal{G}}$, as follows. In this translation from $\mathcal{G}$ to the linear programming instance, the matrix $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a nonzero value (with the exception of self-loops that have only one nonzero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}\cdot x_{j}+a_{ik}\cdot x_{k}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$ and $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), with $a_{ij}=1$, $a_{ik}=-\lambda_{e}$, and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative sign for the minimiser, and the objective function maximises the sum of the values of $x$.
As we use no details of any implementation of a simplex algorithm, we only rehash the principle way it operates here.
From a graphical point of view, the set of constraints, i.e., the rows of $A$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope (including its surface) is a solution.
Broadly speaking, the simplex algorithm traverses the corner of the polytope so that the value the linear objective function takes for the corner is improving – in our case decreasing as we minimise. More precisely, it picks subsets of $n$ inequations that are made sharp: when red as equations, they define a valuation in $\mathbb{R}^{n}$ that satisfies all $m$ inequations; they are bases and define the corners of the simplex.
@@
For non-optimal solutions (w.r.t. the objective function), this is always possible, though in some cases the improvement in (3) is not strict. Even in those cases, there is still a series of changes that all satisfy (1-3) to an optimal solution. Those cases are called degenerate, and while we do not discuss how they are treated in simplex algorithms, we point out that the problems and their solutions are similar to what we describe in Sections 5.2 and 5.3, respectively.
@@ 3. Outline and Motivation Example
In general, the valuation returned as solution is computed by using objective functions based on strategies that are not necessarily optimal. We now adjust the example so that choosing both self-loops is not only non-optimal, but also does not define the correct valuation of the DPG. To this end, we use different discount factors [^5] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1: we choose $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three resulting inequations are (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$; (2) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(b)$; and (3) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(a)$. Choosing both self loops results in the objective function to minimise $1-(\frac{1}{3})\mathsf{val}(a)+\frac{2}{3}\mathsf{val}(b)$.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, according to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We refer to this unique set of values as the evaluation of $\mathbf{b}$, denote by $\mathsf{val}_{\mathbf{b}}$. Note that $\mathsf{val}_{\mathbf{b}}$ is in particular a solution of $H$, so that all all $|E|$ inequations are satisfied.
@@
The valuation $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ of the game is the unique solution of $H$ for which, for all vertices $v$, the inequation to (at least) one of the outgoing edges of $v$ is sharp. This case implies that the edges for which they are sharp describe co-optimal strategies. This solution of $H$ is thus the only one that defines strategies for both players, which shows the second claim.
@@ 5.2. No Local Improvements
  • $S_{\mathsf{val}}^{\sigma}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))\}$ as the set of stale edges; naturally, every vertex has at least one outgoing stale edge: the one defined by $\sigma$;
  • $E_{\mathsf{val}}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=0\}$ as the set of edges, for which the inequation for $\mathsf{val}$ is sharp; in particular, all edges in the basis of $H$ that defines $\mathsf{val}$ are sharp (and stale); and
  • $E_{\mathsf{val}}^{\sigma}$ as any set of edges between $E_{\mathsf{val}}$ and $S_{\mathsf{val}}^{\sigma}$ (i.e. $E_{\mathsf{val}}\subseteq E_{\mathsf{val}}^{\sigma}\subseteq S_{\mathsf{val}}^{\sigma}$) such that $E_{\mathsf{val}}^{\sigma}$ contains an outgoing edge for every vertex $v\in V$; we are interested to deal with sets that retain the game property that every vertex has a successor, we can do that by adding (non-sharp) stale edges to $E_{\mathsf{val}}$.
Note that $S_{\mathsf{val}}^{\sigma}$ is such a set, and, therefore, an adequate set is easy to identify. However, we might be interested in keeping the set small and choosing the edges defined by $E_{\mathsf{val}}$ plus one outgoing edge for every vertex $v$ that does not have an outgoing edge in $E_{\mathsf{val}}$. The most natural solution is to choose the edge $(v,\sigma(v))\in E_{\mathsf{val}}^{\sigma}$ defined by $\sigma$ for each such vertex $v$.
@@
As there is no local improvement, we have $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))\geq\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for every vertex $v$. At the same time, as $(v,\sigma^{\prime}(v))\in E_{\mathsf{val}}^{\sigma}$, we have $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ or $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=0$; together, this provides $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))=\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$. As this holds for every vertex, this provides $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})$. Thus, we now have $f_{\sigma^{\prime}}(\mathsf{val})=f_{\sigma}(\mathsf{val})>0$ and $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})=0$. For an optimal solution $\mathsf{val}^{\prime\prime}$ of $H$ for the objective $f_{\sigma^{\prime}}$, this provides $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})\leq f_{\sigma^{\prime}}(\mathsf{val}_{\varepsilon})<f_{\sigma}(\mathsf{val})$.
@@
When using the most natural choice, $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$, this translates into keeping all transitions, for which the offset is not 0, while changing some of those for which the offset already is 0. This is a slightly surprising choice, since to progress one has to improve on the transitions whose offset is positive, and ignore those with offset 0.

### v5

@@ 2. Preliminaries
The value of a minimiser strategy $\sigma$, denoted by $\mathsf{minival}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{minival}_{\sigma}:v\mapsto\max_{\sigma_{\max}\in\Sigma_{\max}}\mathsf{out}(\rho(v,\sigma,\sigma_{\max}))\ ,$
The value of a maximiser strategy $\sigma$, denoted by $\mathsf{maxival}_{\sigma}:V\rightarrow\mathbb{R}$, is defined as
$\mathsf{maxival}_{\sigma}:v\mapsto\min_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{out}(\rho(v,\sigma_{\min},\sigma))\ .$
@@
We have
$\mathsf{val}_{G}(v)=\max_{\sigma_{\max}\in\Sigma_{\max}}\mathsf{maxival}_{\sigma_{\max}}(v)=\min_{\sigma_{\min}\in\Sigma_{\min}}\mathsf{minival}_{\sigma_{\min}}(v)$
for all $v\in V$. A positional maximiser (resp. minimiser) strategy $\sigma$ is said to be optimal if, and only if, $\mathsf{val}_{\mathcal{G}}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\mathcal{G}}(\sigma(v))$ holds for all maximiser (resp. minimiser) vertices. Such optimal strategies always exist.
The valuation $\mathsf{val}_{\mathcal{G}}$ is the sole valuation that satisfies the fixed point equations
$\mathsf{val}_{\mathcal{G}}(v)=\max_{(v,v^{\prime})\in E}w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}_{\mathcal{G}}(v^{\prime})$
for all maximiser vertices $v\in V_{\max}$ and
$\mathsf{val}_{\mathcal{G}}(v)=\min_{(v,v^{\prime})\in E}w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}_{\mathcal{G}}(v^{\prime})$
for all minimiser vertices $v\in V_{\min}$.
Likewise, we define the value of a pair of strategies $\sigma_{\min}$ and $\sigma_{\max}$, denoted $\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:V\rightarrow\mathbb{R}$, as
$\mathsf{val}_{\sigma_{\min},\sigma_{\max}}:v\mapsto\mathsf{out}(\rho(v,\sigma_{\min},\sigma_{\max}))\ .$
@@
If both of these strategies are optimal, we call $\sigma$ a joint co-optimal strategy. This is the case if, and only if, $\mathsf{val}_{\mathcal{G}}=\mathsf{val}_{\sigma}$ holds [^1] If $\mathsf{val}_{\sigma}=\mathsf{val}_{\mathcal{G}}$ holds, then, for every vertex $v$ and for $v^{\prime}=\sigma(v)$, $\mathsf{val}_{\sigma}(v)=w_{(v,v^{\prime})}+\lambda_{(v,v^{\prime})}\mathsf{val}_{\sigma}(v^{\prime})$ holds. .
@@
For co-optimal strategies $\sigma$, $\mathsf{val}_{\sigma}$ that is equal to $\mathsf{val}_{\mathcal{G}}$ is clearly a solution, as all of these inequations are satisfied, while for a pair of strategies $\sigma$ that is not co-optimal, $\mathsf{val}_{\sigma}$ is not a solution [^3] For every joint strategy $\sigma$ of both players and every vertex $v\in V$, we have $\mathsf{val}_{\sigma}(v)=w_{(v,\sigma(v))}+\lambda_{(v,\sigma(v))}\mathsf{val}_{\sigma}(\sigma(v))$ by definition. Thus, if $\mathsf{val}_{\sigma}$ is a solution, then $\mathsf{val}_{\mathcal{G}}$ also satisfies the fixed point equations that define $\mathsf{val}_{\mathcal{G}}$, and we have that $\mathsf{val}_{\sigma}=\mathsf{val}_{\mathcal{G}}$ holds and that $\sigma$ is co-optimal. .
@@ 2.1. Simplex method
A linear programming problem requires to find a (solution) vector $x\in\mathbb{R}^{n}$ that minimises (or maximises) $c\cdot x$, where $c\in\mathbb{R}^{n}$ is a vector of constants, such that $A\cdot x\leq b$ (or $\geq b$), where $A\in\mathbb{R}^{m\times n}$ is a coefficient matrix and $b\in\mathbb{R}^{m}$ is a vector of constants. Given a game $\mathcal{G}$, we can iteratively define an objective function represented by the vector c, the input matrix $A$, and the vector $b$, such that, in the sequence of linear programming instances, the solution $x$ converges to $\mathsf{val}_{\mathcal{G}}$. In this translation from $\mathcal{G}$ to the linear programming instance, the matrix $A$ is composed of $m$ rows and $n$ columns where $m$ is the number of edges and $n$ the number of vertices in $\mathcal{G}$. Since the constrains in $A$ represent the edges, that are binary relations, only two columns in each row contain a nonzero value (with the exception of self-loops that have only one nonzero coefficient). The vector $b$ contains the weight of the edges. Hence, every row $i$ can be expressed as $a_{ij}\cdot x_{j}+a_{ik}\cdot x_{k}\geq b_{i}$, if $e=(j,k)$ is an edge of $\mathcal{G}$ and $j$ is a maximiser vertex (otherwise the constraint is $\leq b_{i}$), with $a_{ij}=1$, $a_{ik}=-\lambda_{e}$, and $b_{i}=w_{e}$. Finally, $c$ is the unit vector with a positive sign for the maximiser vertices and negative sign for the minimiser, and the objective function maximises the sum of the values of $x$.
As we use no details of any implementation of a simplex algorithm, we only rehash the principle way it operates here. From a graphical point of view, the set of constraints, i.e., the rows of $A$ and $b$, define a region called polytope (or: simplex; hence the name of the approach) in the space $\mathbb{R}^{n}$. Any point that lies within the polytope (including its surface) is a solution.
Broadly speaking, the simplex algorithm traverses the corner of the polytope so that the value the linear objective function takes for the corner is improving – in our case decreasing as we minimise. More precisely, it picks subsets of $n$ inequations that are made sharp: when read as equations, they define a valuation in $\mathbb{R}^{n}$ that satisfies all $m$ inequations; they are bases and define the corners of the simplex.
@@
For non-optimal solutions (w.r.t. the objective function), this is always possible, though in some cases the improvement in (3) is not strict. Even in those cases, there is still a series of changes that all satisfy (1-3) to an optimal solution. Those cases are called degenerate, that is, when the current solution satisfies more inequations than necessary and, although there exists a neighbouring solution leading to a strict improvement, picking it requires a basis change. While we do not discuss how they are treated in simplex algorithms, we point out that the problems and their solutions are similar to what we describe in Sections 5.2 and 5.3, respectively.
@@ 3. Outline and Motivation Example
In general, the valuation returned as solution is computed by using objective functions based on strategies that are not necessarily optimal. We now adjust the example so that choosing both self-loops is not only non-optimal, but also does not define the correct valuation of the DPG. To this end, we use different discount factors [^5] Note that we can also replace the transitions with a smaller discount factor by multiple transitions with a larger discount factor. This would allow for keeping the discount factor uniform, but needlessly complicate the discussion and inflate the size of the example. for the game of Figure 1: we choose $\frac{1}{3}$ for the self-loop of the maximiser vertex and $\frac{2}{3}$ for the other two transitions, so that the three resulting inequations are (1) $\mathsf{val}(a)\leq 1+\frac{2}{3}\mathsf{val}(a)$; (2) $\mathsf{val}(b)\geq\frac{1}{3}\mathsf{val}(b)$; and (3) $\mathsf{val}(b)\geq\frac{2}{3}\mathsf{val}(a)$. Choosing both self loops results in the objective function to minimise $1-\frac{1}{3}\mathsf{val}(a)+\frac{2}{3}\mathsf{val}(b)$.
@@ 5 $\mathsf{val}\leftarrow{}\mathsf{LinearProgramming}(H,f_{\sigma})$
We put a further restriction on $\mathsf{val}$, returned by the call $\mathsf{LinearProgramming}$, in that we require it to be the solution to a basis $\mathbf{b}$ in $H$. Such a basis consists of $|V|$ inequations that are satisfied sharply (again, as equations), such that, according to the definition of basis, these $|V|$ equations uniquely define the values of all vertices. We refer to this unique set of values as the evaluation of $\mathbf{b}$, denote by $\mathsf{val}_{\mathbf{b}}$. Note that $\mathsf{val}_{\mathbf{b}}$ is in particular a solution of $H$, so that all $|E|$ inequations are satisfied, condition that in general does not holds.
@@
The valuation $\mathsf{val}=\mathsf{val}_{\mathcal{G}}$ of the game is the unique solution of $H$ for which, for all vertices $v$, the inequation to (at least) one of the outgoing edges of $v$ is sharp. This case implies that the edges for which they are sharp describe co-optimal strategies. This solution of $H$ is thus the only one that defines strategies for both players, which shows the second claim. The uniqueness of the solution is a consequence of the sharpness: while inequalities define the solution region, equations select exactly one.
@@ 5.2. No Local Improvements
  • $S_{\mathsf{val}}^{\sigma}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))\leq\mathsf{offset}(\mathsf{val},(v,\sigma(v)))\}$ as the set of at least stale edges; naturally, every vertex has at least one outgoing stale edge: the one defined by $\sigma$;
  • $E_{\mathsf{val}}=\{(v,v^{\prime})\in E\mid\mathsf{offset}(\mathsf{val},(v,v^{\prime}))=0\}$ as the set of edges, for which the inequation for $\mathsf{val}$ is sharp; in particular, all edges in the basis of $H$ that defines $\mathsf{val}$ are sharp (and at least stale); and
  • $E_{\mathsf{val}}^{\sigma}$ as any set of edges between $E_{\mathsf{val}}$ and $S_{\mathsf{val}}^{\sigma}$ (i.e. $E_{\mathsf{val}}\subseteq E_{\mathsf{val}}^{\sigma}\subseteq S_{\mathsf{val}}^{\sigma}$) such that $E_{\mathsf{val}}^{\sigma}$ contains an outgoing edge for every vertex $v\in V$; we are interested to deal with sets that retain the game property that every vertex has a successor, we can do that by adding (non-sharp) at least stale edges to $E_{\mathsf{val}}$.
Note that $S_{\mathsf{val}}^{\sigma}$ is such a set, and, therefore, an adequate set is easy to identify. However, we might be interested in keeping the set small and choosing the edges defined by $E_{\mathsf{val}}$ plus one outgoing edge for every vertex $v$ that does not have an outgoing edge in $E_{\mathsf{val}}$. Where there is no local improvement, the most natural solution is to choose the edge $(v,\sigma(v))\in E_{\mathsf{val}}^{\sigma}$ defined by $\sigma$ for each such vertex $v$.
@@
We have $\mathsf{offset}(\mathsf{val},(v,\sigma^{\prime}(v)))\leq\mathsf{offset}(\mathsf{val},(v,\sigma(v)))$ for every vertex $v$, as $\sigma^{\prime}$ only uses the at least stale edges from $E^{\sigma}_{\mathsf{val}}$, so that in particular $f_{\sigma^{\prime}}(\mathsf{val})\leq f_{\sigma}(\mathsf{val})$ holds. We also have $f_{\sigma}(\mathsf{val})>0$ by assumption and $f_{\sigma^{\prime}}(\mathsf{val}^{\prime})=0$ by the optimality of $\mathsf{val}^{\prime}$ for $H^{\prime}$.
For an optimal solution $\mathsf{val}^{\prime\prime}$ of $H$ for the objective $f_{\sigma^{\prime}}$, this provides $f_{\sigma^{\prime}}(\mathsf{val}^{\prime\prime})\leq f_{\sigma^{\prime}}(\mathsf{val}_{\varepsilon})<f_{\sigma}(\mathsf{val})$.
@@
Where $\mathsf{val}$ does not define a strategies of both players, there are always at least $|V|$ inequations sharp, and thus there are vertices with multiple outgoing edges in $E^{\sigma}_{\mathsf{val}}$ that all have 0 offset.
@@
When there is no local improvement, the most natural choice is $E_{\mathsf{val}}^{\sigma}=E_{\mathsf{val}}\cup\{(v,\sigma(v))\mid v\in V\}$: this translates into keeping all transitions, for which the offset is not 0. As a result, $\sigma^{\prime}$ would change some of those for which the offset already is 0, but agree with $\sigma$ for all vertices where this is currently not the case. This is a slightly surprising choice, since to progress one intuitively has to improve on the transitions whose offset is positive—which are then the ones one keeps.